• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name pking

    Log entry time 19:51:45 on April16,2001

    Entry number 60667

    keyword=Accelerator update.

    At about 19:20, J. Grames gave us an update.

    There are some problems with the magnet power supplies for the arc 1 magnets
    and for the hall C extraction magnet. These are expected to take a few hours to
    fix.

    So, there may no beam in this shift. Expect another update toward the end of the
    shift.